# Characterization of the directed landscape from the KPZ fixed point

### v2

@@ 8.3 Monotonicity in general AEPs
The basic coupling of $(\eta^{-}_{t})_{t\geq 0}$ and $(\eta^{+}_{t})_{t\geq 0}$ in Lemma 8.9 can be encoded by a multi-type exclusion process $(\gamma_{t})_{t\geq 0}$, where for each $t\geq 0$ and $x\in\mathbb{Z}$, we let
@@
We interpret the values $1$, 00, $2$, $3$ as representing an (ordinary) particle, a hole, a type-2 particle’, and a type-3 particle’. Each particle, type-2 particle, and type-3 particle jumps according to $p$, whenever the target location is empty. In addition, for any $x\in\mathbb{Z}$ and $v\in\mathbb{Z}\setminus\{0\}$ with $p(v)>0$,
@@
Below we use the notion of a move to denote one operation under these dynamics, i.e. a jump, a swap, or an annihilation. We denote $r=\max\{|v|:p(v)>0\}$ for the maximum range of movement and let $R=10r^{r}$.
55-5- 511-1- 122223333444488881010101022-2- 2002222333355-5- 511-1- 122223333444488881010101022-2- 20022223333
55-5- 511-1- 12222333344441010101022-2- 200333355-5- 511-1- 122223333444488881010101022-2- 20022223333
Figure 2: Top panel: an illustration of the evolution of labels under one jump of a type-2 particle. Bottom panel: an illustration of the evolution of labels under one annihilation, where we randomly choose a type-2 label ($8$) and a type-3 label ($2$) to be removed. (In both panels, black/red/blue/white balls denote ordinary particles/type-2 particles/type-3 particles/holes.)
@@
These rules completely determine the evolution of the labels under either a jump or swap (see the top panel of Figure 2). However, when an annihilation happens, we have some freedom in choosing which pair of labels (one type-2 and one type-3) is removed. We make this choice randomly. Whenever an annihilation happens from a particle jumping from site $x$ to $x+v$, we consider all labels in $\llbracket x-R,x+R\rrbracket$ before the annihilation. Note that there is at least one type-2 label and at least one type-3 label among this set. We uniformly choose one type-2 label to remove, and uniformly choose one type-3 label to remove. The remaining type-2 (resp. type-3) labels are assigned to the remaining type-2 (resp. type-3) particles in the window $\llbracket x-R,x+R\rrbracket$, in the unique way that ensures the labels increase from left to right. (See the bottom panel of Figure 2.)
@@
There are constants $\hat{C}>0$ and $10<\theta<$ 1 depending only on $p$, such that for any type-2 label $X$ and $m\in\mathbb{N}$, the probability that at least $m$ takeovers of $X$ happen (before $X$ is removed in an annihilation) is at most $\hat{C}\theta^{m}$.
@@
There is a constant $10<\vartheta<$ 1 depending only on $p$ such that the following is true. Take any $s\geq 0$. Conditional on the configuration $\gamma_{s}$, if there are sites $x,y\in\mathbb{Z}$ with $\gamma_{s}(x)=2,\gamma_{s}(y)=3$ and $R-2r\leq|x-y|<R-r$, with probability $>\vartheta$ the type-2 label at $x$ will be removed before another $R$ takeovers of it.
@@
Under the setup stated, without loss of generality, we assume that $x<y$, i.e. $y-x\in\llbracket R-2r,R-r-1\rrbracket$. The main task of this proof is to establish the following combinatorial statement.
@@
Claim: Starting from the configuration $\gamma_{s}$, one can construct a sequence of moves within $\llbracket x,y\rrbracket$ (i.e. all jumps, swaps, and annihilations involve a pair of sites in $\llbracket x,y\rrbracket$) such that the last move is an annihilation.
To prove the claim, we introduce some notations. Denote $R_{1}=R/10=r^{r}$. Denote $g_{-}\in\mathbb{N}$ to be the greatest common divisor of $\{v\in\mathbb{Z}_{-}:p(v)>0\}$, and $g_{+}\in\mathbb{N}$ to be the greatest common divisor of $\{v\in\mathbb{Z}_{+}:p(v)>0\}$. In the case where either of the sets is empty, set $g_{-}=0$ or $g_{+}=0$. From the assumption that $\{v:p(v)+p(-v)>0\}$ additively generates $\mathbb{Z}$, $\gcd(g_{-},g_{+})=1$. Note that at least one of $g_{-}$ and $g_{+}$ is nonzero. Without loss of generality we assume that $g_{-}\neq 0$. We let $G_{-}\subset\mathbb{Z}_{-}$ be the semi-group that is additively generated by $\{v\in\mathbb{Z}_{-}:p(v)>0\}$, and $G_{+}\subset\mathbb{Z}_{+}$ be the semi-group that is additively generated by $\{v\in\mathbb{Z}_{+}:p(v)>0\}$. Then we must have $g_{-}\mathbb{Z}\cap(-\infty,-R_{1}]\subset G_{-}$, and $g_{+}\mathbb{Z}\cap[R_{1},\infty)\subset G_{+}$ if $g_{+}\neq 0$.
We now prove the above claim.
Proof: We argue by contradiction. Namely, below we assume that, starting from $\gamma_{s}$, we cannot find a sequence of moves within $\llbracket x,y\rrbracket$ that contains an annihilation.
For any configuration $\gamma:\mathbb{Z}\to\{0,1,2,3\}$, we define its weight to be the sum $\sum_{z\in\llbracket x+1,y-1\rrbracket:\gamma(z)\in\{1,3\}}z$. We now consider all configurations that can be obtained from $\gamma_{s}$ by taking a sequence of moves within $\llbracket x+1,y-1\rrbracket$, and let $\gamma_{*}$ be the one with the smallest weight. If there are multiple such configurations, choose one arbitrarily. Note that by our assumption above, no move in this sequence (from $\gamma_{s}$ to $\gamma_{*}$) can be an annihilation. We then denote $\Gamma=\{z\in\llbracket x+1,y-1\rrbracket:\gamma_{*}(z)\in\{1,3\}\}$. Then for any $z\in\Gamma$, and $v\in\mathbb{Z}_{-}$ with $z+v\geq x+1$ and $p(v)>0$, necessarily $z+v\in\Gamma$. More generally, $(\Gamma+G_{-})\cap\llbracket x+1,y-1\rrbracket\subset\Gamma$.
Now we consider two scenarios: either (1) $\llbracket y-3R_{1},y-1\rrbracket\cap\Gamma=\emptyset$, or (2) there exists some $z_{*}\in\llbracket y-3R_{1},y-1\rrbracket$ with $z_{*}\in\Gamma$.
In the first scenario, starting from $\gamma_{*}$, we can take a sequence of jumps of the type-3 particle that is at $y$ in $\gamma_{*}$, so that it ends up at some $y_{1}\in(y+G_{-})\cap\llbracket y-3R_{1},y-1\rrbracket$. Then if $g_{+}\neq 0$, we can take another sequence of jumps of this type-3 particle, so that it ends up at some $y_{2}\in(y_{1}+G_{+})\cap\llbracket y_{1},y-1\rrbracket$. If $g_{+}=0$, we just let $y_{2}=y_{1}$. Since $g_{-}\mathbb{Z}\cap(-\infty,-R_{1}]\subset G_{-}$, and $g_{+}\mathbb{Z}\cap[R_{1},\infty)\subset G_{+}$ if $g_{+}\neq 0$, and the greatest common divisor of $g_{-}$ and $g_{+}$ is $1$, $y_{2}$ can be any number in $\llbracket y-R_{1}-r,y-R_{1}\rrbracket$. Take any $v_{-}\in\mathbb{Z}_{-}$ such that $p(v_{-})>0$. (As we assumed that $g_{-}\neq 0$, such a $v_{-}$ exists). We can choose $y_{2}$ such that $y_{2}\equiv x\mod|v_{-}|$. Now that there is a type-2 particle at $y_{2}$, and a type-3 particle at $x$, we can take a sequence of moves within $(x+|v_{-}|\mathbb{Z})\cap\llbracket x,y_{2}\rrbracket$, each move involving two locations $z$ and $z+v_{-}$ for some $z\in(x+|v_{-}|\mathbb{Z})\cap\llbracket x+|v_{-}|,y_{2}\rrbracket$, moving all ordinary particles as far as possible to the left and empty sites as far as possible to the right. Then there will be some $z$ and $z+v_{-}$ in $(x+|v_{-}|\mathbb{Z})\cap\llbracket x,y_{2}\rrbracket$, such that there is a type-2 particle at one of these sites and a type-3 particle at the other. Moving one of these particles onto the other then yields an annihilation, which contradicts the assumption.
In the second scenario, we have that $\llbracket x+1,y-4R_{1}\rrbracket\cap(z_{*}+b_{-}\mathbb{Z})\subset\Gamma$, since $\llbracket x+1,y-4R_{1}\rrbracket\cap(z_{*}+g_{-}\mathbb{Z})\subset z_{*}+G_{-}$. If $g_{+}=0$, necessarily $g_{-}=1$, then we have $\llbracket x+1,y-4R_{1}\rrbracket\subset\Gamma$. If $g_{+}\neq 0$, we take any $v_{+}\in\mathbb{Z}_{+}$ such that $p(v_{+})>0$. Then for any $z^{\prime}\in(z_{*}+G_{-})\cap\llbracket x+1,z_{*}-v_{+}\rrbracket$, we must have $z^{\prime}+v_{+}\in\Gamma$. Indeed, otherwise, we can find a sequence of moves leading to a new configuration $\gamma_{*}^{\prime}$, such that for the set
$\Gamma^{\prime}=\{z\in\llbracket x+1,y-1\rrbracket:\gamma_{*}^{\prime}(z)\in\{1,3\}\},$
$\Gamma^{\prime}\setminus\Gamma=\{z^{\prime}+v\}$ and $\Gamma\setminus\Gamma^{\prime}=\{z_{*}\}$. To see that such a sequence exists, let $z_{0}=z_{*},z_{1},\dots,z_{k}=z^{\prime}$ be a sequence with $z_{i}-z_{i+1}\in\{v\in\mathbb{Z}_{-}:p(v)>0\}$. Since $(\Gamma+G_{-})\cap\llbracket x+1,y-1\rrbracket\subset\Gamma$, $z_{i}\in\Gamma$ for all $i$. Then first moving the particle at site $z_{k}$ to $z^{\prime}+v$, then the particle at $z_{k-1}$ to $z_{k}$, $z_{k-2}$ to $z_{k-1}$ and so on until $z_{0}$ moves to $z_{1}$ yields $\gamma_{*}^{\prime}$. The new configuration $\gamma_{*}^{\prime}$ would have a smaller weight than $\gamma_{*}$, contradicting our choice of $\gamma_{*}$.
Therefore, for any $z^{\prime}\in(z_{*}+G_{-})\cap\llbracket x+1,z_{*}-v_{+}\rrbracket$, we must have $(z^{\prime}+G_{+})\cap\llbracket x+1,z_{*}\rrbracket\subset\Gamma$. Combining this with the fact that $\llbracket x+1,y-4R_{1}\rrbracket\cap(z_{*}+g_{-}\mathbb{Z})\subset z_{*}+G_{-}$, and that $\gcd(g_{-},g_{+})=1$, we necessarily have $\llbracket x+2R_{1},z_{*}\rrbracket\subset\Gamma$. Further using that $(z+G_{-})\cap\llbracket x+1,y-1\rrbracket\subset\Gamma$ for any $z\in\Gamma$, we have that $\llbracket x+1,z_{*}\rrbracket\subset\Gamma$. In summary, whether $g_{+}=0$ or not, we have $\llbracket x+1,y-4R_{1}\rrbracket\subset\Gamma$.
At this point we apply a reflection to the entire picture, by defining a new configuration $\gamma^{*}$ as follows: for any $z\in\mathbb{Z}$, we let $\gamma^{*}(z)=1$ if $\gamma_{*}(x+y-z)=0$, $\gamma^{*}(z)=0$ if $\gamma_{*}(x+y-z)=1$, $\gamma^{*}(z)=2$ if $\gamma_{*}(x+y-z)=3$, and $\gamma^{*}(z)=3$ if $\gamma_{*}(x+y-z)=2$. The configuration $\gamma^{*}$ puts us back in scenario (1), and any sequence of moves starting from $\gamma_{*}$ can be reflected to a sequence of moves starting from $\gamma^{*}$. As we showed above, we can create an annihilation starting from $\gamma^{*}$, so we can also find a sequence of moves that reach an annihilation from $\gamma_{*}$. This contradicts our ‘no annihilation’ assumption in the second scenario as well. $\blacksquare$
@@
Now, the number of moves required by the claim is upper bounded by a constant depending only on $p$, since there are only finitely many configurations in $\llbracket x,y\rrbracket$. Therefore, with positive probability depending only on $p$, this exact sequence of moves occurs, and no other move in $\llbracket x-r,y+r\rrbracket$ happens before the end of this sequence. When any annihilation happens (in this sequence of moves), with probability at least $1/(2R)$ the type-2 label at $x$ in $\gamma_{s}$ is removed. Also note that at most $R$ takeovers of the type-2 label happen before the end of this sequence, and so the conclusion follows.
@@
For a type-2 label, we say that it is ready-to-remove if there exists one potential annihilation which could remove it. Then there exists some $10<c_{0}<$ 1 depending only on $p$, such that for any $t>0$, given the configuration at time $t$ with a type-2 label that is ready-to-remove, with probability at least $c_{0}$ the type-2 label will be removed before its next takeover.
Below we take a type-2 label $X$ at time 00. Let $t_{0}=0$. For any $i\in\mathbb{N}$, let $t_{i}$ be the time of its $i$-th takeover, and set $t_{i}=\infty$ if less than $i$ takeovers happen to label $X$. For $i\in\mathbb{Z}_{\geq 0}$, we then let $\tau_{i}$ be the first time after $t_{i}$, such that either $X$ is ready-to-remove, or there is a type-3 particle to its right, with distance $\geq R-2r$ and $<R-r$. Then $\tau_{i}$ is a stopping time, and may equal $\infty$. (In particular, if $t_{i}=\infty$ we have $\tau_{i}=\infty$.)
For any $i\in\mathbb{Z}_{\geq 0}$, we note that at most $R$ takeovers can happen to $X$ between $t_{i}$ and $\tau_{i}$. Also given $\tau_{i}$ which is <absent<$\infty$< ∞, by Lemma 8.13 and the statement at the beginning of this proof, with probability $>\vartheta\wedge c_{0}$ the label $X$ will be removed before $R$ takeovers after $\tau_{i}$. Therefore, we have that $\mathbb{P}[t_{i+R}<\infty\mid t_{i}<\infty]<1-\vartheta\wedge c_{0}$. Thus the conclusion follows. ∎
@@
Let $C_{*}$ be a sufficiently large constant depending on $p$. Consider the following events:
@@
    Every type-2 or type-3 label that enters $\llbracket-w-3R,w+3R\rrbracket$ between time 00 and $t$ is contained in the set $\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$ at time 00.
@@
    For every type-2 or type-3 label that is contained in $\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$ at time 00, at most $m$ takeovers happen (before it is removed).
Claim: Under $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, we must have $Y_{t,w}\leq m+2R$.
Proof: For any $x\in\llbracket-w,w\rrbracket$, we consider the last time (before $t$) that a type-2 or type-3 label goes from $\leq x-1$ to $\geq x$, or vice versa. This also includes the event where an annihilation happens between a label $\leq x-1$ and a label $\geq x$. If no such time exists, then $\eta^{-}_{t}(x)-\eta^{+}_{t}(x)=\eta^{-}_{0}(x)-\eta^{+}_{0}(x)\leq 0$. Otherwise, let $s\in(0,t]$ be the last such time. Without loss of generality, we assume that a type-2 label goes from $\leq x-1$ to $\geq x$ at time $s$, or it is at $\leq x-1$ right before time $s$ and annihilates with a type-3 label $\geq x$ at time $s$. Denote the trajectory of that type-2 label by $(w(s^{\prime}))_{0\leq s^{\prime}<s}$. Then we have $\eta^{-}_{t}(x)-\eta^{+}_{t}(x)\leq\eta^{-}_{s-}(w(s-))-\eta^{+}_{s-}(w(s-))+2R$. On the other hand, since $\eta^{-}_{0}(w(0))\leq\eta^{+}_{0}(w(0))$, we have that for any $0\leq s^{\prime}<s$, $\eta^{-}_{s^{\prime}}(w(s^{\prime}))-\eta^{+}_{s^{\prime}}(w(s^{\prime}))$ is at most the number of takeovers that happened to this type-2 label between time 00 and $s^{\prime}$. By $\mathcal{E}_{1}$ we have that $w(0)\in\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$. By $\mathcal{E}_{2}$ we have that at most $m$ takeovers happen to this type-2 label. Thus $\eta^{-}_{s-}(w(s-))-\eta^{+}_{s-}(w(s-))\leq m$, further implying that $\eta^{-}_{t}(x)-\eta^{+}_{t}(x)\leq m+2R$. $\blacksquare$
@@
By Lemma 8.12, we have $\mathbb{P}[\mathcal{E}_{2}]>1-C(t+w)\theta^{m}$. As for $\mathcal{E}_{1}$, we consider any type-2 or type-3 label that is at location $x<-C_{*}(t+w)$ at time 00. We can couple it with a random walk with a different jump rate $p^{\prime}$ (depending on $p$), such that it is always to the left of that random walk. Therefore, by taking $C_{*}$ large enough (depending on $p$), with have that with probability $>1-C\exp(c(w+x))$, it is to the left of $-w-3R$ up to time $t$. By taking a union bound over $x<-C_{*}(t+w)$, we have that with probability $>1-C\exp(-c(t+w))$, any type-2 or type-3 label that is to the left of $-C_{*}(t+w)$ at time 00 is to the left of $-w-3R$ up to time $t$.
Similarly, we have that with probability $>1-C\exp(-c(t+w))$, any type-2 or type-3 label that is to the right of $C_{*}(t+w)$ at time 00 is to the right of $w+3R$ up to time $t$. Thus we get that $\mathbb{P}[\mathcal{E}_{1}]>1-C\exp(-ct)$, and so $\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}]>1-C(t+w)\theta^{m}-C\exp(-ct)$. The conclusion follows. ∎

### v3

@@ 7.3 Monotonicity in general AEPs
Throughout this subsection, we consider AEPs with the jump distribution $p$, such that $p(1)>1/2$, $p(i)=0$ for all $|i|>K$, and $|p(i)|\leq K$ for all $i\leq K$, for some $K\in\mathbb{N}$. We use $C,c>0$ to denote constants depending only on $K$, and their values may change from line to line.
The basic coupling of $(\eta^{-}_{t})_{t\geq 0}$ and $(\eta^{+}_{t})_{t\geq 0}$ (with height functions $(h^{-}_{t})_{t\geq 0}$ and $(h^{+}_{t})_{t\geq 0}$) in Lemma 7.9 can be encoded by a multi-type exclusion process $(\gamma_{t})_{t\geq 0}$, where for each $t\geq 0$ and $x\in\mathbb{Z}$, we let
@@
We interpret the values $1$, 0, 22, 33 as representing an (ordinary) particle, a hole, a type-2 particle’, and a type-3 particle’. Each particle, type-2 particle, and type-3 particle jumps according to $p$, whenever the target location is empty. In addition, for any $x\in\mathbb{Z}$ and $v\in\mathbb{Z}\setminus\{0\}$ with $p(v)>0$,
@@
Below we use the notion of a move to denote one operation under these dynamics, i.e. a jump, a swap, or an annihilation.
$5-51-12233448810102-202233$
5-51-122334410102-20335-51-12233448810102-202233
Figure 2: Top panel: an illustration of the evolution of labels under one jump of a type-2 particle. Bottom panel: an illustration of the evolution of labels under one annihilation, where we randomly choose a type-2 label (88) and a type-3 label (22) to be removed. (In both panels, black/red/blue/white balls denote ordinary particles/type-2 particles/type-3 particles/holes.)
@@
These rules completely determine the evolution of the labels under either a jump or swap (see the top panel of Figure 2). However, when an annihilation happens, we have some freedom in choosing which pair of labels (one type-2 and one type-3) is removed. We make this choice randomly. Whenever an annihilation happens from a particle jumping from site $x$ to $x+v$, we consider all labels in $\llbracket x-K,x+K\rrbracket$ before the annihilation. Note that there is at least one type-2 label and at least one type-3 label among this set. We uniformly choose one type-2 label to remove, and uniformly choose one type-3 label to remove. The remaining type-2 (resp. type-3) labels are assigned to the remaining type-2 (resp. type-3) particles in the window $\llbracket x-K,x+K\rrbracket$, in the unique way that ensures the labels increase from left to right. (See the bottom panel of Figure 2.)
@@
There are constants $\hat{C}>0$ and $0<\theta<1$ depending only on $K$, such that for any type-2 label $X$ and $m\in\mathbb{N}$, the probability that at least $m$ takeovers of $X$ happen (before $X$ is removed in an annihilation) is at most $\hat{C}\theta^{m}$.
@@
There is a constant $0<\vartheta<1$ depending only on $K$ such that the following is true. Take any $s\geq 0$. Conditional on the configuration $\gamma_{s}$, if there are sites $x,y\in\mathbb{Z}$ with $\gamma_{s}(x)=2,\gamma_{s}(y)=3$ and $|x-y|\leq K$, with probability $\vartheta$ the type-2 label at $x$ will be removed before another takeover of it.
@@
Under the setup stated, without loss of generality, we assume that $x<y$, i.e. $y-x\in\llbracket 1,K\rrbracket$. We have the following combinatorial statement.
@@
Claim: Starting from the configuration $\gamma_{s}$, one can construct a sequence of moves within $\llbracket x,y\rrbracket$, such that each move is one of
  • (1)
    an ordinary, or type-2, or type-3 particle jumps from $z$ to $z+1$, for some $z\in\llbracket x,y-1\rrbracket$;
  • (2)
    a particle at $z$ swaps with a type-2 or type-3 particle at $z+1$, for some $z\in\llbracket x,y-1\rrbracket$;
  • (3)
    a type-2 (resp. type-3) particle at $z$ annihilates with a type-3 (resp. type-2) particle at $z+1$, for some $z\in\llbracket x,y-1\rrbracket$.
Besides, the last move of this sequence is an annihilation.
Proof: Via jumps and swaps as stated in (1) and (2), all the particles are moved to the right and all the holes are moved to the left. Namely, for some $x^{\prime},y^{\prime}\in\llbracket x,y\rrbracket$, there is a hole at every integer in $\llbracket x,x^{\prime}-1\rrbracket$, and a particle at every integer in $\llbracket y^{\prime}+1,y\rrbracket$. For each $z\in\llbracket x^{\prime},y^{\prime}\rrbracket$, there is a type-2 or type-3 particle at $x$. Note that since $\gamma_{s}(x)=2$ and $\gamma_{s}(y)=3$, after such moves there is at least one type-2 particle and one type-3 particle in $\llbracket x^{\prime},y^{\prime}\rrbracket$. Then in particular, there is a type-2 particle next to a type-3 particle in $\llbracket x^{\prime},y^{\prime}\rrbracket$, and one can take an annihilation between them. $\blacksquare$
@@
Now, the number of moves required by the claim is upper bounded by a constant depending only on $K$, since there are only finitely many configurations in $\llbracket x,y\rrbracket$. Therefore, using that $p(1)>1/2$, with positive probability depending only on $K$, this exact sequence of moves occurs, and no other move in $\llbracket x-K,y+K\rrbracket$ happens before the end of this sequence. When any annihilation happens (in this sequence of moves), with probability at least $1/(2K)$ the type-2 label at $x$ in $\gamma_{s}$ is removed. Also, note that no takeover of the type-2 label happens before the end of this sequence, and so the conclusion follows.∎
@@
Below we take a type-2 label $X$ at time 0. Let $t_{0}=0$. For any $i\in\mathbb{N}$, let $t_{i}$ be the time of its $i$-th takeover, and set $t_{i}=\infty$ if less than $i$ takeovers happen to label $X$. For $i\in\mathbb{Z}_{\geq 0}$, we then let $\tau_{i}$ be the first time $\geq t_{i}$, such that there is a type-3 particle to the right of $X$, with distance $\leq K$. Then $\tau_{i}$ is a stopping time, and may equal $\infty$. (In particular, if $t_{i}=\infty$ we have $\tau_{i}=\infty$.)
For any $i\in\mathbb{Z}_{\geq 0}$, we note that no takeover can happen to $X$ between $t_{i}$ and $\tau_{i}$. This is because immediately before any takeover of $X$, there must be a type-3 particle to the right of $X$ and with distance $\leq K$. Also given $\tau_{i}$ which is $\infty$, by Lemma 7.13, with probability $\vartheta$ the label $X$ will be removed before any takeover after $\tau_{i}$. Therefore, we have that $\mathbb{P}[t_{i+1}<\infty\mid t_{i}<\infty]<1-\vartheta$. Thus the conclusion follows. ∎
@@
Let $C_{*}$ be a sufficiently large constant depending on $K$. Consider the following events:
@@
    Every type-2 or type-3 label that enters $\llbracket-w-K,w+K\rrbracket$ between time 0 and $t$ is contained in the set $\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$ at time 0.
@@
    For every type-2 or type-3 label that is contained in $\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$ at time 0, at most $m/2$ takeovers happen (before it is removed).
Claim: Under $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, we must have $Y_{t,w}\leq m+2K$.
Proof: For any $x\in\llbracket-w,w\rrbracket$, we consider the last time (before $t$) that a type-2 or type-3 label goes from $\leq x-1$ to $\geq x$, or vice versa. This also includes the event where an annihilation happens between a label $\leq x-1$ and a label $\geq x$. If no such time exists, then $h^{-}_{t}(x)-h^{+}_{t}(x)=h^{-}_{0}(x)-h^{+}_{0}(x)\leq 0$. Otherwise, let $s\in(0,t]$ be the last such time. Without loss of generality, we assume that a type-2 label goes from $\leq x-1$ to $\geq x$ at time $s$, or it is at $\leq x-1$ right before time $s$ and annihilates with a type-3 label $\geq x$ at time $s$. Denote the trajectory of that type-2 label by $(w(s^{\prime}))_{0\leq s^{\prime}<s}$. Then we have $h^{-}_{t}(x)-h^{+}_{t}(x)\leq h^{-}_{s-}(w(s-))-h^{+}_{s-}(w(s-))+2K$.
On the other hand, since $h^{-}_{0}(w(0))\leq h^{+}_{0}(w(0))$, we have that for any $0\leq s^{\prime}<s$, $(h^{-}_{s^{\prime}}(w(s^{\prime}))-h^{+}_{s^{\prime}}(w(s^{\prime})))/2$ is at most the number of takeovers that happened to this type-2 label between time 0 and $s^{\prime}$. By $\mathcal{E}_{1}$ we have that $w(0)\in\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$. By $\mathcal{E}_{2}$ we have that at most $m$ takeovers happen to this type-2 label. Thus $h^{-}_{s-}(w(s-))-h^{+}_{s-}(w(s-))\leq m$, further implying that $h^{-}_{t}(x)-h^{+}_{t}(x)\leq m+2K$. $\blacksquare$
@@
By Lemma 7.12, we have $\mathbb{P}[\mathcal{E}_{2}]>1-C(t+w)\theta^{m/2}$. As for $\mathcal{E}_{1}$, we consider any type-2 or type-3 label that is at location $x<-C_{*}(t+w)$ at time 0. We can couple it with a random walk that jumps to the right by distance $K$, with rate $K^{2}$, such that the label is always to the left of that random walk. Therefore, by taking $C_{*}$ large enough (depending on $K$), we have that with probability $1-C\exp(c(w+x))$, it is to the left of wK-w-K up to time $t$. By taking a union bound over $x<-C_{*}(t+w)$, we have that with probability $1-C\exp(-c(t+w))$, any type-2 or type-3 label that is to the left of $C_{*}(t+w)$ at time 0 is to the left of wK-w-K up to time $t$.
Similarly, we have that with probability $1-C\exp(-c(t+w))$, any type-2 or type-3 label that is to the right of $C_{*}(t+w)$ at time 0 is to the right of $w+K$ up to time $t$. Thus we get that $\mathbb{P}[\mathcal{E}_{1}]>1-C\exp(-ct)$, and so $\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}]>1-C(t+w)\theta^{m/2}-C\exp(-ct)$. The conclusion follows. ∎